Use the method of substitution to solve the system of linear equations.
3x-y=5
7x-2y=13

Answer by reviewermath(1029)

3x - y = 5 ---> y = 3x - 5
7x - 2y = 13 ---> 7x - 2(3x - 5) = 13
7x - 6x + 10 = 13
x = 3
y = 3(3) - 5 = 4
Answer: x = 3 and y = 4